calculate the ratio of diffusion rates for carbon monoxide and carbon dioxide


1
Expert's answer
2020-12-03T14:07:42-0500

V1/V2 = √(Mr2/Mr1) according to Graham's law of effusion (or diffusion)

V(CO)/V(CO2) = √((44 g/mol)/(28 g/mol)) = 1.2536